OpenStudy (anonymous):

I need help and an explaination on how to solve: 3x/5 - 4/5(x=1) = 2 - 3/10(3x-4)

OpenStudy (anonymous):

can't really make sense of your equation. Use equation editor or draw in the board

OpenStudy (anonymous):

you have two equal signs in the equation

OpenStudy (anonymous):

\[3x/5 - 4/(5 ) (x+1)=2- 3/10(3x-4)\]

OpenStudy (anonymous):

maybe \[\frac{3x}{5}-\frac{4}{5}(x+1)=2-\frac{3}{10}(3x-4)\]?

OpenStudy (anonymous):

if so, multiply both sides by 10 to get \[6x-8(x+1)=20-3(3x-4)\] and go from there

OpenStudy (anonymous):

if not, ignore what i wrote

OpenStudy (anonymous):

prodgers2, I suggest that you go to the following link if you are using windows. http://www.microsoft.com/download/en/details.aspx?id=15702 It is a FREE math program by Microsoft that can factor and solve equations for you and give you all the steps used to solve that equation. It is great for homework and other things. You will not regret it because it is an awesome tool to help reinforce your math.